Exactly how many cubic centimeters of ice cream can an ice cream cone hold (within the cone) if its height is 8 centimeters and the radius of its base is 3 centimeters? Include correct units with your solution. Use "pi" for π.
Found 2 solutions by lwsshak3, wilft1:
Answer by lwsshak3(11628)

Exactly how many cubic centimeters of ice cream can an ice cream cone hold (within the cone) if its height is 8 centimeters and the radius of its base is 3 centimeters? Include correct units with your solution. Use "pi" for π.
**
Volume of cone=(1/3)πr^2h
=(π*3^2*8)/3=72π/3=24π
ans:
Ice cream cone can hold 24π cubic cm of ice cream

Answer by wilft1(217)

the formula for the volume of a cone is pi r squared, times height, divide by 3
soooooo...

3.14 x 9 x 8 = 226.08 cu/cm divide by 3, = 75.36 cu/cm